# Vista-LLaMA: Reliable Video Narrator via Equal Distance to Visual Tokens

### v1

@@ 1 Introduction
Figure 1: Video language processing with LLaMA [24] and our Vista-LLaMA. The vanilla LLaMA treats visual tokens (Refer to caption) the same as other language tokens (Refer to caption), weakening the impact for tokens in long distance. Our model retains the same mechanism for language tokens and strengthens the impact of the visual tokens. The intensity of the impact of each token is conveyed through the depth of the line color (Refer to caption). Our model provides the accurate response for the presented scenario.
The surge in multi-modal vision-and-language models [1, 12, 14], capable of comprehending both visual (e.g., image/video) and language data, can be attributed to the recent achievements of large language models (LLMs) such as GPT [8], GLM [35], and LLaMA [37]. Video-language models (video-LMs) pose greater challenges in scaling due to increased computational and annotation costs compared to image-language models (image-LMs). Recent research has therefore focused on effectively training video-LMs by utilizing pre-trained image-LMs [19, 37].
Video-LMs benefit from this warm-start approach to enhance visual representation learning by projecting video frames into language space and treating videos as several prompt language tokens [32, 15]. However, this diminishes the visual impact of videos on text generation and lacks explicit temporal modeling in videos. The generated text is often not related to the video content, as depicted in Figure 1, a phenomenon known as hallucination in language processing  [6]. The distance between the generated text token and the visual tokens in large language models may be a contributing factor, especially when the visual tokens are distant from the generated text. Additionally, handling longer videos poses a challenge because of the constraints on context length in large language models, where numerous visual tokens take up a substantial portion of the context.
@@
We demonstrate the effectiveness of Vista-LLaMA  on four challenging video question answering (QA) benchmarks, where our method outperforms several previous works, and achieves the state-of-the-art in zero-shot NExT-QA [28], and MSRVTT [30]. We also show that our attention design and the temporal modeling mechanism improves capacity of large language model on video-question answering by a large margin. Comprehensive experiments are conducted to demonstrate the effectiveness of our designs. We summarize our contributions as follows:
@@ 2.1 Large Language Models
Large language models (LLMs) [26, 39, 16, 2, 25] have demonstrated exceptional proficiency in understanding language and reasoning, resulting in the generation of high-quality natural language text in diverse domains. LLMs have already initiated a technological revolution and have found extensive application in various domains. Additionally, a series of open source large models, including LLaMA [24], and OPT [38], have significantly contributed to technological advancements and made remarkable contributions to the NLP community. Leveraging the foundation established by these impressive LLMs, researchers have further expanded their capabilities and developed exceptional models for diverse NLP tasks (e.g. Vicuna [4]). Our work is also built upon these remarkable LLMs, and we equip the language models with a novel attention mechanism and a sequential visual projector that enhances their abilities to comprehend visual content in videos.
@@ 2.2 Multi-modal Large Language Models
Researchers have been diligently exploring the application of LLMs for processing multi-modal problems [13, 9]. The existing approaches can be classified into two primary categories. In the first category, LLMs are treated as controllers while the multi-modal models as tools [10]. When presented with the specific task, the LLMs first interact with user instructions and decide which tools to employ. Subsequently, it generates comprehensive responses by amalgamating the outcomes derived from these readily available multi-modal models. These approaches, such as Visual ChatGPT [27], HuggingGPT [18], and ViperGPT [22], have shown impressive results on various multi-modal tasks without training models. For the second category, large-scale multi-modal models are trained on the multi-modal data. The principle behind this line of researches is to align other modal pre-trained models with textual LLMs. Flamingo [1] incorporates a perceiver resampler and a gated cross-attention layer to connect a frozen image encoder with the LLM. BLIP2 [12] introduces a Q-Former to map each image into fixed-length tokens in the language embedding space via the learned queries. LLaVA [14], mPLUG-owl [33], and MiniGPT4 [39] develop the image-LLMs utilizing image-instruction training pairs. Video Chat [13] extends image encoders to enable large models to comprehend visual content in videos. Video-ChatGPT [15] is trained on video instructional data to give appropriate answers for multi-modal inputs. In this work, we improve the understanding of videos in a video-language model by incorporating temporal modeling and a novel multi-modal attention mechanism.
@@ 2.3 Video Question Answering
Video Question Answering (VideoQA) is a task that involves answering language questions that are based on a given video [30]. It requires the ability to understand and reason across different semantic levels, which in turn demands a capacity for multi-modal understanding. Previous VideoQA benchmarks primarily concentrated on short videos, asking questions according to the specified visual facts (e.g., location and objects) [31, 7]. Recently, several new benchmarks have been proposed to focus on resolving temporal and causal questions in longer video clips [23]. NExT-QA [28] is an example of such a benchmark, aiming to uncover the causalities or intentions of particular events, and infer subsequent actions within the video. In this work, we investigate the video-language model on zero-shot open-ended VideoQA where no training question-answer pairs are provided in the training stage.
@@ 3.1 Overview
  • Visual Encoder. We utilize pre-trained EVA-CLIP-g [21] and ViT-L [17] as visual encoder. The last layer of ViT encoder is removed because it specializes in aggregating the features of the first token for contrastive learning.
@@
  • Pre-trained Large Language Model. In this study, we utilize LLaVa [14], which is fine-tuned on Vicuna-7B [4] using instructional image-text pairs, for processing videos and texts. The causal mask is employed in all attention processes, incorporating the attention interplay between visual and text tokens. The pre-trained language model is frozen in the present study, and a new attention mechanism is developed to maintain a consistent distance to visual tokens for all textual tokens.
@@ 3.2 Equal Distance to Visual Tokens
EDVT-Attention. The vanilla attention model lacks positional awareness, with no encoded relative distance for natural language processing. In contrast, Rotary Positional Embeddings (RoPE) [20] encodes the position data of tokens using a rotation matrix, which inherently includes an explicit relative position dependency. Within each attention layer, RoPE is implemented across all projected query and key inputs in order to compute the attention weights via leveraging relative distance between tokens. The query situated at the $j^{th}$ position incorporates rotary position embedding through $\tilde{\mathbf{q}}_{j}\!=\!\mathbf{R}_{j}\mathbf{q}_{j}$, wherein $\mathbf{R}_{j}\!\in\!\mathbb{R}^{d\times d}$ represents the rotary matrix for the $j^{th}$ position. Consequently, the attention involving relative position embedding is expressed as follows:
@@ 3.3 Sequential Visual Projector
The visual projector aims to map video features into the language embedding space, allowing for the fusion and processing of visual and textual inputs by the substantial language model. As shown in Figure 3, earlier visual projectors either employ the linear layer or the query transformer (Q-Former) [12] to directly convert frame features into language tokens. However, the lack of temporal relationship in these methods impedes thorough video understanding in LLMs. We introduce the sequential visual projector in Figure 3 to encode the temporal context into the visual tokens.
@@
Moreover, we can represent the entire video with fewer tokens by merely sampling a small number of visual tokens. This method tackles the challenge of encoding lengthy videos by employing sequential encoding. By integrating prior visual context into subsequent visual tokens, we can accomplish an adequate representation of the entire video through sparse sampling of projected frame features. Besides, this technique permits larger language models to manage much longer videos without constraints of context length.
@@ 4.1 Experimental Setup
  • MSVD-QA [30] is a dataset built upon Microsoft Research Video Description Corpus [3], commonly used in video caption tasks. The MSVD-QA dataset comprises a total of 1,970 video clips with 50,505 QA pairs.
@@
  • Open-Ended Zero-Shot Question-Answer Evaluation. As video-language models generate responses of varying lengths to open-ended questions, it is challenge to evaluate models with traditional word matching strategy. We employ LLM-Assisted evaluation, in line with [15], for assessment purposes. Given the question, correct answer, and predicted answer by model, GPT-3.5 model is used to return True or False judgement and relative score (00 to $5$).
  • Video-Based Text Generation Benchmarking. We evaluate the text generation performance following [15] from five aspects: Correctness of Information, Consistency, Detail Orientation, Contextual Understanding, and Temporal Understanding. The test set for this evaluation is based on the ActivityNet-200 [7], featuring videos with rich, dense descriptive captions and associated question-answer pairs from human annotations. The evaluation pipeline is also built with the GPT-3.5 model and relative score (00 to $5$) is generated.
  Method NExT-QA [28] MSVD-QA [30] MSRVTT-QA [30] ActivityNet-QA [34]
@@
FrozenBiLM [32] - - 32.2 - 16.8 - 24.7 -
Video Chat [13] 56.2 3.2 56.3 2.8 45.0 2.5 26.5 2.2
LLaMA Adapter [37] - - 54.9 3.1 43.8 2.7 34.2 2.7
Video LLaMA [36] - - 51.6 2.5 29.6 1.8 12.4 1.1
MovieChat [19] 49.9 2.7 61.0 2.9 49.7 2.8 51.5 3.1
Video-ChatGPT [15] 54.6 3.2 64.9 3.3 49.3 2.8 35.2 2.7
@@ 4.2 Comparison to State-of-the-Arts
For the zero-shot open-ended video question answering tasks, we compare our model with FrozenBiLM [32], Video Chat [13], LLaMA Adapter [37], VideoLLaMA [36], Video-ChatGPT [15], and MovieChat [19] in Tab. 1. FrozenBiLM adapts frozen the bidirectional language model pre-trained on web-scale multi-modal data, showing promising results in zero-shot VideoQA settings. Other compared models are built on recent large auto-regressive language models. Despite pre-existing models have produced substantial results, Vista-LLaMA  consistently outperforms them, achieving state-of-the-art (SoTA) performance across three datasets: NExT-QA [28], MSVD-QA [30], and MSVTT-QA [30]. Our method obtains the highest results, with 60.5% accuracy on MSRVTT, elevating the performance of the second-best model by nearly 10%. Additionally, our method attains 60.7% accuracy on NExT-QA, markedly superior to Video-ChatGPT [15], which employs similar architectures as ours and is trained using the same datasets. These results demonstrate Vista-LLaMA’s capability to comprehend video content and produce precise answers.
@@
Video Chat [13] 2.23 2.24 2.50 2.53 1.94
LLaMA Adapter [37] 2.03 2.15 2.32 2.30 1.98
Video LLaMA [36] 1.96 1.79 2.18 2.16 1.82
Video-ChatGPT [15] 2.40 2.37 2.52 2.62 1.98
@@
We present the results of the evaluation of video-based text generation in Tab. 2. The results reveal its competent performance across all key aspects when compared with the recently introduced video-language models, Video Chat[13], VideoLLaMA [36], and Video-ChatGPT [15]. Despite being trained with identical datasets, our model outperforms Video-ChatGPT in four aspects. Our mothod offers a more comprehensive interpretation, and its responses are more in tune with the overarching context of the video content than comparable approaches. By utilizing the EDVT-Attention and sequential temporal modeling techniques, our model demonstrates a strong ability to generate text that is contextually appropriate, detailed, and includes precise timing for video inputs.
@@ 4.3 Abalation Study
Effect of Design Choices. We investigate the effect of our two designs, including the equal-distance to visual-tokens (EDVT) attention and the sequential visual projector. As shown in Figure 4, our model is tested on NExT-QA [28], which comprises three different types of questions. The baseline model employs LLaVA [14] as the language model, and Q-Former in BLIP-2 [12] as the visual projector. We first incorporate temporal modeling into the baseline, utilizing the projected tokens from the previous frame as the query tokens to generate visual tokens for the current frame, a setup denoted as “w Seq” in Figure 4. Further enhancing this version with EDVT-Attention, we introduce “w Seq&EDVT”, ultimately forming our final Vista-LLaMA. The results affirm the efficacy of our design across all question types.
Figure 4: Comparison of different design choices on NExT-QA [28]4.3). For clarity, accuracy of base model and two variants are given (“Baseline”: the frozen LLaVA [14] with trainable Q-Former [12], “w Seq”: base model with sequential visual projector, “w Seq&EDVT”: base model with both sequential visual projector and EDVT-Attention).
EDVT-Attention. We here validate the effectiveness of our core EDVT-Attention. Table 3 reports the comparison results of different models in combination with our EDVT-Attention on NExT-QA [28]. Initially, we integrate the EDVT-Attention into Video-ChatGPT [15], where only the attention component is substituted with our EDVT-Attention during training. Further deployment of the EDVT-Attention into Video-ChatGPT [15] yields significant performance gains (e.g., 65.1% escalates to 72.8% for causal questions). The accuracy is enlarged across all three setting types, indicating that our EDVT-Attention considerably enhances the multi-modal understanding in large language models.
@@
Video-ChatGPT [15] 37.6 65.1 54.9 54.6
@@
Table 3: Comparison of EDVT-Attention design with diffent visual encoders on NExT-QA [28]4.3). For clarity, accuracy of all questions and three types of questions are reported (“Tem.”: temporal, “Cau.”: causal, “Des.”: descriptive, “Avg.”: average).
@@
Table 4: Comparison of different visual projectors on NExT-QA [28]. The linear projector is initialized with pre-trained weights in LLaVa [14]. BERT init. and BLIP-2 init. indicate that the visual projector is initialized with weights from BERT [5] and BLIP-2 [12]. SeqQ-Former is the proposed sequential visual projector. See §4.3 for more details.
Visual Projector. The visual projector is trainable within the model. Therefore, we have examined the impacts of different visual projector designs, as shown in Table 4. We evaluated three visual projector variants. The Q-Former, initialized using BLIP-2 [12], considerably outperforms the Q-Former that was initialized with BERT [5]. This indicates that visual projectors, when pre-trained on image-text pairs, can significantly enhance video comprehension. The model’s performance peaks when integrating the sequential design, referred to as “SeqQ-Former” in the figure, outperforming all other visual projectors in terms of accuracy.
Figure 6: The effect of training visual tokens on NExT-QA [28]. Accuracy of all questions and three types of questions, including temporal (Tem.), descriptive (Des.), and causal (Cau.), are presented with different colors.
Figure 7: Visualization results on different video questions. The questions and annotated answers are located on the left side. The generated text from Video-ChatGPT [29] and our model is presented in the green and orange boxes, respectively. See §4.3 for more details.
@@
Quantitative Results. We visualize the generated responses of Video-ChatGPT [15] and Vista-LLaMA  for different videos in Figure 7. The frames sampled at different timestamps is presented, with questions listed below the images. In the first video, the mobile phone is turned, sequentially exposing each side over time. The question asks for the side of the mobile phone at the end of the video. Video-ChatGPT gives the wrong response, as the camera is never present at the bottom left side in the video. The hallucination occurs in Video-ChatGPT, whereas our model predicts the answer correctly. The responses in three cases demonstrate that Video-ChatGPT provides unrelated answers that do not correspond to the video content. In the second case, the person in pink is always standing while the man is sitting on the ground, but Video-ChatGPT incorrectly states that the person in pink is also sitting on the ground. In the third video, the man is consistently walking up the mountain and never walking down. However, Video-ChatGPT falsely claims that the man continues walking down the mountain. It is quite common for the video language model to produce unrelated responses. In contrast, our model provides the correct and reliable responses, aligning with the video content. This demonstrates that our model significantly reduces hallucination in video understanding and delivers more accurate responses.

### v2

# Vista-LLaMA: Reducing Hallucination in Video Language Models via Equal Distance to Visual Tokens

@@ 1 Introduction
Figure 1: Video language processing with LLaMA [Touvron2023LLaMAOA] and our Vista-LLaMA. The vanilla LLaMA treats visual tokens (Refer to caption) the same as other language tokens (Refer to caption), weakening the impact for tokens in long distance. Our model retains the same mechanism for language tokens and strengthens the impact of the visual tokens. The intensity of the impact of each token is conveyed through the depth of the line color (Refer to caption). Our model provides the accurate response for the presented scenario.
The surge in multi-modal vision-and-language models [alayrac2022flamingo, Li2023BLIP2BL, Liu2023VisualIT], capable of comprehending both visual (e.g., image/video) and language data, can be attributed to the recent achievements of large language models (LLMs) such as GPT [floridi2020gpt], GLM [zeng2022glm], and LLaMA [zhang2023llama]. Video-language models (video-LMs) pose greater challenges in scaling due to increased computational and annotation costs compared to image-language models (image-LMs). Recent research has therefore focused on effectively training video-LMs by utilizing pre-trained image-LMs [song2023moviechat, zhang2023llama].
Video-LMs benefit from this warm-start approach to enhance visual representation learning by projecting video frames into language space and treating videos as several prompt language tokens [yang2022zero, maaz2023video]. However, this diminishes the visual impact of videos on text generation and lacks explicit temporal modeling in videos. The generated text is often not related to the video content, as depicted in Figure 1, a phenomenon known as hallucination in language processing  [dhuliawala2023chain]. The distance between the generated text token and the visual tokens in large language models may be a contributing factor, especially when the visual tokens are distant from the generated text. Additionally, handling longer videos poses a challenge because of the constraints on context length in large language models, where numerous visual tokens take up a substantial portion of the context.
@@
We demonstrate the effectiveness of Vista-LLaMA  on four challenging video question answering (QA) benchmarks, where our method outperforms several previous works, and achieves the state-of-the-art in zero-shot NExT-QA [Xiao2021NExTQANP], and MSRVTT [Xu2017VideoQA]. We also show that our attention design and the temporal modeling mechanism improves capacity of large language model on video-question answering by a large margin. Comprehensive experiments are conducted to demonstrate the effectiveness of our designs. We summarize our contributions as follows:
@@ 2.1 Large Language Models
Large language models (LLMs) [Scao2022BLOOMA1, Zhu2023MiniGPT4EV, OpenAI2023GPT4TR, brown2020language, Tsimpoukelli2021MultimodalFL] have demonstrated exceptional proficiency in understanding language and reasoning, resulting in the generation of high-quality natural language text in diverse domains. LLMs have already initiated a technological revolution and have found extensive application in various domains. Additionally, a series of open source large models, including LLaMA [Touvron2023LLaMAOA], and OPT [Zhang2022OPTOP], have significantly contributed to technological advancements and made remarkable contributions to the NLP community. Leveraging the foundation established by these impressive LLMs, researchers have further expanded their capabilities and developed exceptional models for diverse NLP tasks (e.g. Vicuna [chiang2023vicuna]). Our work is also built upon these remarkable LLMs, and we equip the language models with a novel attention mechanism and a sequential visual projector that enhances their abilities to comprehend visual content in videos.
@@ 2.2 Multi-modal Large Language Models
Researchers have been diligently exploring the application of LLMs for processing multi-modal problems [li2023videochat, Gao2022MISTM, ma2023temporal, lu2024zero]. The existing approaches can be classified into two primary categories. In the first category, LLMs are treated as controllers while the multi-modal models as tools [gupta2023visual]. When presented with the specific task, the LLMs first interact with user instructions and decide which tools to employ. Subsequently, it generates comprehensive responses by amalgamating the outcomes derived from these readily available multi-modal models. These approaches, such as Visual ChatGPT [Wu2023VisualCT], HuggingGPT [shen2023hugginggpt], and DoraemonGPT [yang2024doraemongpt], have shown impressive results on various multi-modal tasks without training models. For the second category, large-scale multi-modal models are trained on the multi-modal data. The principle behind this line of researches is to align other modal pre-trained models with textual LLMs. Flamingo [alayrac2022flamingo] incorporates a perceiver resampler and a gated cross-attention layer to connect a frozen image encoder with the LLM. BLIP2 [Li2023BLIP2BL] introduces a Q-Former to map each image into fixed-length tokens in the language embedding space via the learned queries. LLaVA [Liu2023VisualIT], mPLUG-owl [Ye2023mPLUGOwlME], and MiniGPT4 [Zhu2023MiniGPT4EV] develop the image-LLMs utilizing image-instruction training pairs. Video Chat [li2023videochat] extends image encoders to enable large models to comprehend visual content in videos. Video-ChatGPT [maaz2023video] is trained on video instructional data to give appropriate answers for multi-modal inputs. In this work, we improve the understanding of videos in a video-language model by incorporating temporal modeling and a novel multi-modal attention mechanism.
@@ 2.3 Video Question Answering
Video Question Answering (VideoQA) is a task that involves answering language questions that are based on a given video [Xu2017VideoQA]. It requires the ability to understand and reason across different semantic levels, which in turn demands a capacity for multi-modal understanding. Previous VideoQA benchmarks primarily concentrated on short videos, asking questions according to the specified visual facts (e.g., location and objects) [Xu2016MSRVTTAL, ma2022weakly]. Recently, several new benchmarks have been proposed to focus on resolving temporal and causal questions in longer video clips [Tapaswi2015MovieQAUS]. NExT-QA [Xiao2021NExTQANP] is an example of such a benchmark, aiming to uncover the causalities or intentions of particular events, and infer subsequent actions within the video. In this work, we investigate the video-language model on zero-shot open-ended VideoQA where no training question-answer pairs are provided in the training stage.
@@ 3.1 Overview
  • Visual Encoder. We utilize pre-trained EVA-CLIP-g [sun2023eva] and ViT-L [radford2021learning] as visual encoder. The last layer of ViT encoder is removed because it specializes in aggregating the features of the first token for contrastive learning.
@@
  • Pre-trained Large Language Model. In this study, we utilize LLaVa [Liu2023VisualIT], which is fine-tuned on Vicuna-7B [chiang2023vicuna] using instructional image-text pairs, for processing videos and texts. The causal mask is employed in all attention processes, incorporating the attention interplay between visual and text tokens. The pre-trained language model is frozen in the present study, and a new attention mechanism is developed to maintain a consistent distance to visual tokens for all textual tokens.
@@ 3.2 Equal Distance to Visual Tokens
EDVT-Attention. The vanilla attention model lacks positional awareness, with no encoded relative distance for natural language processing. In contrast, Rotary Positional Embeddings (RoPE) [su2021roformer] encodes the position data of tokens using a rotation matrix, which inherently includes an explicit relative position dependency. Within each attention layer, RoPE is implemented across all projected query and key inputs in order to compute the attention weights via leveraging relative distance between tokens. The query situated at the $j^{th}$ position incorporates rotary position embedding through $\tilde{\mathbf{q}}_{j}\!=\!\mathbf{R}_{j}\mathbf{q}_{j}$, wherein $\mathbf{R}_{j}\!\in\!\mathbb{R}^{d\times d}$ represents the rotary matrix for the $j^{th}$ position. Consequently, the attention involving relative position embedding is expressed as follows:
@@ 3.3 Sequential Visual Projector
The visual projector aims to map video features into the language embedding space, allowing for the fusion and processing of visual and textual inputs by the substantial language model. As shown in Figure 3, earlier visual projectors either employ the linear layer or the query transformer (Q-Former) [Li2023BLIP2BL] to directly convert frame features into language tokens. However, the lack of temporal relationship in these methods impedes thorough video understanding in LLMs. We introduce the sequential visual projector in Figure 3 to encode the temporal context into the visual tokens.
@@
Moreover, we can represent the entire video with fewer tokens by merely sampling a small number of visual tokens. This method tackles the challenge of encoding lengthy videos by employing sequential encoding. By integrating prior visual context into subsequent visual tokens, we can accomplish an adequate representation of the entire video through sparse sampling of projected frame features. Besides, this technique permits larger language models to manage much longer videos without length limitation.
@@ 4.1 Experimental Setup
  • MSVD-QA [Xu2017VideoQA] is a dataset built upon Microsoft Research Video Description Corpus [Chen2011CollectingHP], commonly used in video caption tasks. The MSVD-QA dataset comprises a total of 1,970 video clips with 50,505 QA pairs.
@@
  • Open-Ended Zero-Shot Question-Answer Evaluation. As video-language models generate responses of varying lengths to open-ended questions, it is challenge to evaluate models with traditional word matching strategy. We employ LLM-Assisted evaluation, in line with [maaz2023video], for fair comparison. Given the question, correct answer, and predicted answer, GPT3.5-turbo-0613 is used to return True or False judgement and relative score (00 to $5$).
  • Video-Based Text Generation Benchmarking. We evaluate the text generation performance following [maaz2023video] from five aspects: Correctness of Information, Consistency, Detail Orientation, Contextual Understanding, and Temporal Understanding. The test set for this evaluation is based on the ActivityNet-200 [caba2015activitynet], featuring videos with rich, dense descriptive captions and associated question-answer pairs from human annotations. The evaluation pipeline is also built with the GPT-3.5 model and relative score (00 to $5$) is generated.
  Method NExT-QA [Xiao2021NExTQANP] MSVD-QA [Xu2017VideoQA] MSRVTT-QA [Xu2017VideoQA] ActivityNet-QA [Yu2019ActivityNetQAAD]
@@
FrozenBiLM [yang2022zero] - - 32.2 - 16.8 - 24.7 -
Video Chat [li2023videochat] 56.2 3.2 56.3 2.8 45.0 2.5 26.5 2.2
LLaMA Adapter [zhang2023llama] - - 54.9 3.1 43.8 2.7 34.2 2.7
Video LLaMA [zhang2023video] - - 51.6 2.5 29.6 1.8 12.4 1.1
MovieChat [song2023moviechat] 49.9 2.7 61.0 2.9 49.7 2.8 51.5 3.1
Video-ChatGPT [maaz2023video] 54.6 3.2 64.9 3.3 49.3 2.8 35.2 2.7
@@ 4.2 Comparison to State-of-the-Arts
For the zero-shot open-ended video question answering tasks, we compare our model with FrozenBiLM [yang2022zero], Video Chat [li2023videochat], LLaMA Adapter [zhang2023llama], VideoLLaMA [zhang2023video], Video-ChatGPT [maaz2023video], and MovieChat [song2023moviechat] in Tab. 1. FrozenBiLM adapts frozen the bidirectional language model, showing promising results in zero-shot VideoQA settings. Other compared models are built on recent large auto-regressive language models. Despite pre-existing models have produced substantial results, Vista-LLaMA  consistently outperforms them, achieving state-of-the-art (SoTA) performance across three datasets: NExT-QA [Xiao2021NExTQANP], MSVD-QA [Xu2017VideoQA], and MSVTT-QA [Xu2017VideoQA]. Our method obtains the highest results, with 60.5% accuracy on MSRVTT, elevating the performance of the second-best model by nearly 10%. Additionally, our method attains 60.7% accuracy on NExT-QA, markedly superior to Video-ChatGPT [maaz2023video]. These results demonstrate Vista-LLaMA’s capability to comprehend video content and produce precise answers.
@@
Video Chat [li2023videochat] 2.23 2.24 2.50 2.53 1.94
LLaMA Adapter [zhang2023llama] 2.03 2.15 2.32 2.30 1.98
Video LLaMA [zhang2023video] 1.96 1.79 2.18 2.16 1.82
Video-ChatGPT [maaz2023video] 2.40 2.37 2.52 2.62 1.98
@@
We present the results of the evaluation of video-based text generation in Tab. 2. The results reveal its competent performance across all aspects when compared with the recent video-language models, Video Chat[li2023videochat], VideoLLaMA [zhang2023video], and Video-ChatGPT [maaz2023video]. Despite being trained with identical datasets, our model outperforms Video-ChatGPT in four aspects. Our mothod offers a more comprehensive interpretation, and its responses are more in tune with the overarching context of the video content than comparable approaches. By utilizing the EDVT-Attention and sequential temporal modeling techniques, our model demonstrates a strong ability to generate text that is contextually appropriate, detailed, and includes precise timing for video inputs.
@@ 4.3 Abalation Study
Effect of Design Choices. We investigate the effect of our two designs, including the equal-distance to visual-tokens (EDVT) attention and the sequential visual projector. As shown in Figure 4, our model is tested on NExT-QA [Xiao2021NExTQANP], which comprises three different types of questions. The baseline model employs LLaVA [Liu2023VisualIT] as the language model, and Q-Former in BLIP-2 [Li2023BLIP2BL] as the visual projector. We first incorporate temporal modeling into the baseline, utilizing the projected tokens from the previous frame as the query tokens to generate visual tokens for the current frame, a setup denoted as “w Seq” in Figure 4. Further enhancing this version with EDVT-Attention, we introduce “w Seq&EDVT”, ultimately forming our final Vista-LLaMA. The results affirm the efficacy of our design across all question types.
Figure 4: Comparison of different design choices on NExT-QA [Xiao2021NExTQANP]4.3). For clarity, accuracy of base model and two variants are given (“Baseline”: the frozen LLaVA [Liu2023VisualIT] with trainable Q-Former [Li2023BLIP2BL], “w Seq”: base model with sequential visual projector, “w Seq&EDVT”: base model with both sequential visual projector and EDVT-Attention).
EDVT-Attention. We here validate the effectiveness of our core EDVT-Attention. Table 3 reports the comparison results of different models in combination with our EDVT-Attention on NExT-QA [Xiao2021NExTQANP]. Initially, we integrate the EDVT-Attention into Video-ChatGPT [maaz2023video], where only the attention component is substituted with our EDVT-Attention during training. Further deployment of the EDVT-Attention into Video-ChatGPT [maaz2023video] yields significant performance gains (e.g., 65.1% escalates to 72.8% for causal questions). The accuracy is enlarged across all three setting types, indicating that our EDVT-Attention considerably enhances the multi-modal understanding in large language models.
@@
Video-ChatGPT [maaz2023video] 37.6 65.1 54.9 54.6
@@
Table 3: Comparison of EDVT-Attention design with diffent visual encoders on NExT-QA [Xiao2021NExTQANP]4.3). For clarity, accuracy of all questions and three types of questions are reported (“Tem.”: temporal, “Cau.”: causal, “Des.”: descriptive, “Avg.”: average).
@@
Table 4: Comparison of different visual projectors on NExT-QA [Xiao2021NExTQANP]. The linear projector is initialized with pre-trained weights in LLaVa [Liu2023VisualIT]. BERT init. and BLIP-2 init. indicate that the visual projector is initialized with weights from BERT [devlin2018bert] and BLIP-2 [Li2023BLIP2BL]. SeqQ-Former is the proposed sequential visual projector. See §4.3 for more details.
Visual Projector. The visual projector is trainable within the model. Therefore, we have examined the impacts of different visual projector designs, as shown in Table 4. We evaluated three visual projector variants. The Q-Former, initialized using BLIP-2 [Li2023BLIP2BL], considerably outperforms the Q-Former that was initialized with BERT [devlin2018bert]. This indicates that visual projectors, when pre-trained on image-text pairs, can significantly enhance video comprehension. The model’s performance peaks when integrating the sequential design, referred to as “SeqQ-Former” in the figure, outperforming all other visual projectors in terms of accuracy.
Figure 6: The effect of training visual tokens on NExT-QA [Xiao2021NExTQANP]. Accuracy of all questions and three types of questions, including temporal (Tem.), descriptive (Des.), and causal (Cau.), are presented with different colors.
Figure 7: Visualization results on different video questions. The questions and annotated answers are located on the left side. The generated text from Video-ChatGPT [Xu2023BaizeAO] and our model is presented in the green and orange boxes, respectively. See §4.3 for more details.
@@
Quantitative Results. We visualize the generated responses of Video-ChatGPT [maaz2023video] and Vista-LLaMA  for different videos in Figure 7. The frames sampled at different timestamps is presented, with questions listed below the images. In the first video, the mobile phone is turned, sequentially exposing each side over time. The question asks for the side of the mobile phone at the end of the video. Video-ChatGPT gives the wrong response, as the camera is never present at the bottom left side in the video. The hallucination occurs in Video-ChatGPT, whereas our model predicts the answer correctly. The responses in three cases demonstrate that Video-ChatGPT provides unrelated answers that do not correspond to the video content. In the second case, the person in pink is always standing while the man is sitting on the ground, but Video-ChatGPT incorrectly states that the person in pink is also sitting on the ground. In the third video, the man is consistently walking up the mountain and never walking down. However, Video-ChatGPT falsely outputs that the man continues walking down the mountain. In contrast, our model provides the correct and reliable responses. This demonstrates that our model significantly reduces hallucination in video understanding and delivers more accurate responses.